OpenStudy (anonymous):

simplify log(9x^5)+5log(1/x)

ganeshie8 (ganeshie8):

log(9x^5)+5log(1/x) = 5log(9x) + 5log1 - 5logx =5[log(9x/x)] =5log9 =10log3